InfoMall World Wide Web Services

The World Wide Web

Commercial use of the Internet, and particularly the World Wide Web, is growing at an incredible rate. Companies are realizing that they can very inexpensively advertise their products and services to a rapidly-growing, global population of Internet users that already numbers in the millions.

InfoMall Web Services

The Northeast Parallel Architectures Center (NPAC) at Syracuse University specializes in both the high-performance computing and communications technologies that are the basis of the World Wide Web infrastructure, and the digital multimedia and information technologies that are the basis of Web applications.

The InfoMall technology transfer program at NPAC offers a wide range of World Wide Web services aimed at providing New York State companies with a strong corporate presence on the Internet. These services range from the development of a corporate “home page” on the World Wide Web to advanced client/server applications using the Web. This enables corporations to provide on-line advertising, marketing and information services.

InfoMall offers consulting and development services for:

• Hypertext document creation

• Applications interfaces using forms and CGI programs

• Digital multimedia, including video-on-demand

• Web interfaces to database and other client/server applications

• Strategies for marketing using the Web

• Electronic commerce

• Virtual reality and interactive applications using VRML and Java

• Building World Wide Web servers

InfoMall also allows you to serve your documents from our high-end Web server running on a powerful multi-processor Silicon Graphics PowerChallenge computer with a highspeed Internet connection. The server can also be accessed via the NYNET gigabit network testbed, which uses leading-edge fiber-optic ATM networking technologies to provide a prototype for the Information Superhighway of the future. You can find the Info-Mall World Wide Web server by directing your Web client (e.g. Mosaic or Netscape) to

http://www.infomall.org/.

InfoMall Web services start at $5000 per year, which includes creation of a basic corporate home page on the InfoMall Web server, and consulting on how the Web can be used to benefit your company. More advanced services are also available at an additional fee.